OpenStudy (tomektomeknyc):

Hello, how to find 4 sin 5 Pi/12

OpenStudy (solomonzelman):

pi= 180

OpenStudy (tomektomeknyc):

can you solve it completely?

OpenStudy (cwrw238):

i guess you want to find this value without using a calculator?

OpenStudy (tomektomeknyc):

yes

OpenStudy (cwrw238):

ok pi/12 = 180/12 = 15 degrees one way would be to use the relation sin(a - b) = sina cosb - cosa sinb let a = 60 and b = 45

OpenStudy (cwrw238):

you can use the exact values of sines cosines of 45 and 60 degrees from the standard triangles: